OpenStudy (candy13106):

Identify the slope of a line that would parallel to the given line. 1. -4x+2y=4 2. y-5=3(x-2)

OpenStudy (here_to_help15):

long time huh :P

OpenStudy (candy13106):

yep

OpenStudy (here_to_help15):

mkay lemme see

OpenStudy (here_to_help15):

Im stumped :/ @mathstudent55 @iGreen

OpenStudy (candy13106):

lol

OpenStudy (here_to_help15):

lol :-/

OpenStudy (here_to_help15):

@paki

OpenStudy (here_to_help15):

well i g2g ;/ but cya around next time maybe i will be able to help

OpenStudy (igreen):

Which line is it??

OpenStudy (candy13106):

ok

OpenStudy (igreen):

The first one or the second one?

OpenStudy (candy13106):

there is no line idk y my teacher gave this homework to me

OpenStudy (here_to_help15):

Your help is here :D @campbell_st

OpenStudy (candy13106):

both

OpenStudy (igreen):

-4x+2y=4 Change to Slope-Intercept Form(y = mx + b): -4x + 2y = 4 Add 4x to both sides: 2y = 4x + 4 Divide 2 to all terms: y = 2x + 2 Now it's in the form of y = mx + b, where m = slope. So what's the slope of the y = 2x + 2? @Candy13106

OpenStudy (candy13106):

slope=2

OpenStudy (igreen):

Yep, and Parallel lines have the same slope..so the slope of the second line will be 2.

OpenStudy (igreen):

That's for the first equation.

OpenStudy (candy13106):

ok wat about the second question

OpenStudy (igreen):

y - 5 = 3(x - 2) is in the form of y1 - y = m(x - x1), where m = slope, so what's the slope of y - 5 = 3(x - 2)? @Candy13106

OpenStudy (campbell_st):

point slope form \[y = y_{1} = m(x - x_{1})~~~~~point~~(x_{1}, y_{1})~~~~ slope = m\] what is the slope in the 2nd line

OpenStudy (candy13106):

slope=3

OpenStudy (igreen):

Yep, since parallel lines have the same slope, the slope of the 2nd line for y - 5 = 3(x - 2) is 3.

OpenStudy (candy13106):

oh ok thanks for helping

OpenStudy (igreen):

-4x + 2y = 4 Slope of 2nd line = 2 y - 5 = 3(x - 2) Slope of 2nd line = 3

OpenStudy (igreen):

No problem! @Candy13106

OpenStudy (anonymous):

I green can u help me

OpenStudy (igreen):

Sure just make your own question and tag me using the @ key.

OpenStudy (anonymous):

A. |–6|, 0, 2, –|–5| B. –|5|, |–4|, 2, 5 C. –9, –|–2|, 0, |9| D. |–10|, –|–4|, 1, |–7|